AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 06:33:40 2025
Author
*
:
Subject
*
:
<p>01:17</p> <p>Run 40 continues.</p> <p>Beam intensity looks OK. Oscillates between just below 2.5e7 and just below 3e7. (attachment 1)</p> <p>Target density 4.3e11, YR09 pressure 1.54e-10 mbar.</p> <p>Rates in monitoring code show beam on target.</p> <p>FEE temperatures all OK - plot over time (attachment 2)</p> <p>System check OK</p> <p>Leakage current OK -trending down (attachment 3)</p> <p>Tape server and merger OK. Merger stats (attachment 4)</p> <p>ADC data items looks OK, spectrum browser rates OK</p> <p>Spectrum browser stat (attachment 7 LOG scale) - looks like some HEC were firing previously. Spec browser histograms zero'd to monitor</p> <p>pn + nn waveforms (attachments 5+6)</p> <p>Monitoring code energy histograms (attachment 8)</p> <p>02:40</p> <p>X2go screen size has decreased further -> cannot see target density but can view beam intensity. Will use ML vac plot to help gauge beam on target</p> <p>Beam intensity has decreased over the last hour but has now stabilised and is now ~2e7 at injection. Pressure in YR09 has decreased slightly accordingly from ML online but not significantly (still >1e-10 mbar)</p> <p>Monitoring code rate plots indicate beam on target</p> <p>All daq checks OK</p> <p>Will continue to monitor beam intensity</p> <p>04:14</p> <p>Ring intensity has increased and appears stable back at 2.5e7.</p> <p>Target density has slightly decreased to 3e11, ring pressure YR09 stable at 1.5e-10 mbar. I am reluctant to try and move the nozzle while my x2go screen is not 100%.</p> <p>FEE temps OK</p> <p>System checks OK</p> <p>Merger and tape server OK</p> <p>Rates in all FEEs look OK</p> <p>Spectrum browser stats (attchment 9) no HEC firing</p> <p>Monitoring code rates look OK -> beam on target (attachment 10)</p> <p>05:14</p> <p>Beam intensity looks OK, Target density 3.11e11. Ring pressure YR09 2e-10 mbar.</p> <p>All checks OK</p> <p>Monitoring code rate looks OK -> beam on target</p> <p>06:20</p> <p>Beam intensity OK, Target density 3.1e11. Pressure YR09 1.9e-10 mbar.</p> <p>FEE temps OK</p> <p>System checks OK</p> <p>Statistics and rates of all FEEs OK</p> <p>Merger and Tape server OK</p> <p>Monitoring code rates look OK -> beam on target</p> <p>Merger statistics (attachment 11)</p> <p>7:15</p> <p>Beam intensity 2E7, target density getting bad 2eE11. WIll fix it from the lab.</p> <p>FEE temps OK</p> <p>System checks OK</p> <p>Merger OK</p> <p>Stats OK</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown